Test marks (%) obtained in a test are normally distributed with a mean of 55 % and variance of 400 (% squared).

what is the probability that a randomly selected student scored below 40%?


1
Expert's answer
2021-05-30T20:41:05-0400

"P(X<40)=P(Z<\frac{40-55}{20})=P(Z<-0.75)=0.2266."